If 4.86 g of solid carbon dioxide was placed into an evacuated 2.00 L container, sealed, and then heated until it completely changed to a gas with a final temperature of 75 °C, what would be the pressure?
1
Expert's answer
2012-10-30T10:01:13-0400
pV = nRT
m(CO2) = 4.86 g
V = 2 L
t = 75 deg C
--------------------
p - ?

Solution:

n(CO2) = m(CO2) / M(CO2) = 4.86 g / 44 g/mol = 0.110 molT = t + 273 = 348 K R = 0.082 L*atm/(K*mol) p = nRT / V = 0.082 L*atm/(K*mol) *
0.11 mol * 348 K / 2 L = 1.57 atm

Answer: p = 1.57 atm